Benitez Company currently outsources a relay switch that is a component in one of its products. The switches cost $23 each. The company is considering making the switches internally at the following projected annual production costs:

Unit-level material cost $ 4
Unit-level labor cost $ 3
Unit-level overhead $ 2
Batch-level set-up cost (6,000 units per batch) $ 32,000
Product-level supervisory salaries $ 41,000
Allocated facility-level costs $ 27,000

The company expects an annual need for 6,000 switches. If the company makes the product, it will have to utilize factory space currently being leased to another company for $2,200 a month. If the company decides to make the parts, total costs will be:

Multiple Choice

  • $25,600 less than if the switches are purchased.

  • $15,400 more than if the switches are purchased.

  • $42,400 more than if the switches are purchased.

  • $27,000 less than if the switches are purchased.
Answer #1

Answer C

$42,400 more than if the switches are purchased.

From the above problem variable cost per switch = material + labor + overhead = 4+3+2 = $9 per unit

Cost Benefit Analysis :

Cost for 6000 Switches if made by company :

Variable Cost         = 6000 * 9 = $54,000

Fixed Cost = 32,000+41,000+27,000 = $100,000

Lost lease rent per year = 2,200 * 12 = $26,400

Total Costs = $180,400

Savings in purchase costs = 6000 * $23 = $138,000

Net increase in costs = 180,400 - 138,000 = $42,400.